Jocelynn:

Juan drove 126 miles in 3 hours. If he continued at the same rate, how long would it take to travel 336 miles?

snowflake0531:

Write a proportion\[\frac{ 126 }{ 3 } = \frac{ 336 }{ x } \\ 126x = 3(336)\] you can solve from there